# A Buddy for Betelgeuse: Binarity as the Origin of the Long Secondary Period in Orionis

### v1

@@ 4.8 Hypothesis 8: Modulation by a companion
In this section, we overview the ways in which a companion could induce an LSP. This family of solutions is the generally favored mechanism for producing LSPs, as this class of scenarios naturally explains the long timescale and steady period as the orbital time of the companion. As discussed in SectionĀ 2, the existence of a P-L relation for LSPs is not an intrinsic problem, as orbital architecture is sensitive to the stellar mass and radius. A remaining point of tension is the relatively high fraction of long-period variables with LSPs and RV modulations compared to main-sequence stars (Nicholls etĀ al., 2009), which can be addressed by appealing to mass growth of the orbiting body (e.g. Livio & Soker 1984; see discussion in §5 of SoszyÅski etĀ al. 2021). At least regarding RSGs, the observed binary fraction is very high for massive stars (Sana etĀ al., 2012; de Mink etĀ al., 2013), and this fraction does not include undetected lower-mass companions (we return to this in SectionĀ 5.2).
@@ 4.8.3 Hypothesis 8c: Circumstellar Modulation Near Companion
| Misidentified FM | 4.1 | ā | X | ? | - | ? | ā |
| Giant convective cells | 4.2 | ā? | ? | ā? | - | - | X |
@@
| Non-radial pulsation | 4.6 | ā | ā? | ? | ? | ? | X |
| Dust- | 4.7 | ā | ā | X | ā | ā | ā? |
| Binarity: tidal | 4.8.1 | X | ā | ? | ? | ? | ā |
| Binarity: occultation | 4.8.2 | ā | ā | ā | ā | X | ā |
| Binarity: dust modulation | 4.8.3 | ā | ā | ā | ā | ā | ā |
@@
We thus conclude that circumstellar dust modulation by a companion is the most plausible explanation for the LSP in Betelgeuse. Within this framework, we can use the observational data available to estimate the properties of a companion consistent with the RV and RV-LC phase difference. This companion is hereafter referred to as 222āBā stands for Betelbuddy. The properties (SectionĀ 5) and detectability (SectionĀ 5.2) of are discussed subsequently.
@@ 4.9 Summary of Hypotheses
TableĀ 1 provides a quick-look reference for the explanatory power and limitations of each theory discussed in this section.
@@ 5.1 Properties of Companion
where is the velocity amplitude and d is the LSP period from the RV observations of Granzer etĀ al. (2022), and is the present-day mass estimate of from Joyce etĀ al. (2020). For simplicity, we assume a circular orbit. This gives us a lower mass limit of . This value is an order of magnitude larger than the mass estimates calculated for LSPs in RGBs with RV constraints, which are typically in the or less range, i.e., closer to brown dwarfs than to dwarf stars (Nicholls etĀ al., 2009). In order to cause substantial modulation over the course of an orbit, we assume that a relatively low inclination is needed for any interaction with the circumstellar dust to be observable from our vantage point. Therefore we estimate that the mass of is unlikely to exceed .
@@
which recovers an orbital separation of assuming the present-day mass of Betelgeuse estimated by Joyce etĀ al. (2020). Using their stellar radius of yields an orbital distance of . This is well outside of the CO2-H2O MOLsphere described by MontargĆØs etĀ al. (2021), which extends to . It also places the L1 Lagrange point at , or . In this orbital configuration, the surface of Betelgeuse is not expected to deform significantly, as the Roche radius for this orbit is , and the equipotential corresponding to is nearly spherical.
@@ 5.2 Observability of Companion
Even at its highest plausibleāsay āit would be nearly impossible to detect close to such a luminous and intrinsically variable star (the average brightness of Betelgeuse hovers near ) except in some exotic and improbable scenarios (see SectionĀ 5.2.1). Even adopting an optimistic 10 for , the luminosity contrast would still be , likely undetectable in integrated light. This challenge is exacerbated by the fact that Betelgeuseās intrinsic variability from convection and pulsations entails mag fluctuations in the V-band (see our Fig.2), with intrinsic fluctuations on the order of more than , at least 1000 times the intrinsic brightness of .
Spectral differencing at different phases would be another possible avenue for detection, but again, subtle changes in the of Betelgeuse and changes on the FM timescale would overwhelm variations in spectra unless the companion were substantially hotter than Betelgeuse. At its hottest, a 2 star might be spectral class F, corresponding to a of K. A companion of this temperature would only just barely be visible on the Wien tail in the UV (a 7000K blackbody outshines a brighter, 3600K blackbody only above Hz, or below ), and realistically, this is a drastic overestimate of the temperature of .
@@ 5.2.1 Exotic Scenarios
While the companionās mass and proximity to a massive star might lead one to speculate that it may be a neutron star or other exotic compact body, as far as the authors of this study are aware, no such detections have been made (Posson-Brown etĀ al., 2006; Kashyap etĀ al., 2020). Future X-ray observations scheduled at favorable epochs might rule this out conclusively. Such an experiment might also help characterize the presence of a possible corona around a non-compact companion, which might yield insight into the exact mechanism for interaction with any circumstellar dust.
@@ 5.2.2 Observability Under Normal Conditions
Another possibility is to inspect the shape of evolving asymmetries in sub-milimeter observations of the dusty region around Betelgeuse using an instrument like ALMA. There is indeed substantial asymmetry in the dust observed near the radius of the predicted orbital separation of Richards etĀ al. (e.g. 2013); OāGorman etĀ al. (e.g. 2017, 2020); Kervella etĀ al. (e.g. 2018b).
@@
A similar observing strategy was undertaken by OāGorman etĀ al. (2015), who took VLA measurements at 0.7, 1.3, 2.0, 3.5, and 6.1cm at multiple different phase points along the LSP over a 4-year cycle (in addition to single-epoch observations by Lim etĀ al. (1998) 4 years prior). Their Fig.Ā 1 does indicate some time-dependent modulation of an asymmetric circumstellar envelope visible in the 0.7cm interferometric data around mas333At maximum separation, will be at mas for a stellar optical photosphere diameter of 43 mas), but it is far from conclusive evidence, as an 8-year time span entails measurements during multiple different LSP cycles (orbits of ). While we do expect intrinsic diversity in dust structures on a replenishment timescale commensurate with the FM as the FM interacts with the large-scale convection, would cause a systematic modulation in the circumstellar dust (the proposed mechanism for the LSP itself).
@@ 6 Conclusions
In Section 5, we calculated physical properties of the companion and the Orionis orbital system as constrained by exiting observations. A summary of these parameters is provided in TableĀ 2. We discussed why has not been detected before, commented on exotic physical scenarios that would be theoretically observable, discussed practical limitations on the companionās detectability, and described the advances in method, technique and instrumentation needed to detect it in the future.
@@
The presence of dense circumstellar material is of reinvigorated interest after the recent Supernova 2023ixf, the closest core-collapse event of the decade, which showed strong signatures of early interaction in an asymmetric circumstellar medium (e.g. Hiramatsu etĀ al., 2023; Jacobson-GalĆ”n etĀ al., 2023; Hosseinzadeh etĀ al., 2023; Smith etĀ al., 2023; Teja etĀ al., 2023; Zimmerman etĀ al., 2024; Martinez etĀ al., 2024; Singh etĀ al., 2024; Moriya & Singh, 2024; Bostroem etĀ al., 2024) around a pulsating Red Supergiant progenitor (e.g. Soraisam etĀ al., 2023; Jencson etĀ al., 2023; Kilpatrick etĀ al., 2023). Moreover, supernova remnants are now being observed with unprecedented resolution in the era of JWST (e.g. recent work by Temim etĀ al., 2024; Milisavljevic etĀ al., 2024), and the proposed observations of Betelgeuseās surrounding environment, whether or not they support the presence of , will be valuable in constraining the expected level and geometry of circumstellar asymmetries that may imprint on the remnant morphology (Polin etĀ al., 2022; Mandal etĀ al., 2023, 2024). If the existence of is confirmed and a progenitor identified, one could also predict its own future evolution in the context of Betelgeuseās eventual explosion. Such simulations have recently been carried out in the case of Type Ia SNe with a surviving companion (Bauer etĀ al., 2019; Bhat etĀ al., 2024; Wong etĀ al., 2024).
Observations of Betelgeuse have identified photometric variations on a 30 year period consistent with a rotational km/s (Joyce etĀ al., 2020), which for an inclination angle of 20ā recovers an equatorial velocity of 15Ā km/s (see Dupree etĀ al. 1987; Gilliland & Dupree 1996; Uitenbroek etĀ al. 1998; Kervella etĀ al. 2009, 2018b, and review byWheeler & Chatzopoulos 2023). This has led to speculation that Betelgeuse may be a merger product (e.g. Wheeler etĀ al., 2017; Chatzopoulos etĀ al., 2020; Shiber etĀ al., 2024). This hypothesis was invoked to account for fast envelope rotation which would not the survive spin-down as Betelgeuse ascended up the RSG branch in a single-star evolutionary scenario. Although the large-scale convective envelope itself might be the source of this RV signal (Ma etĀ al., 2024), if the apparent RV dipole signal persists at a consistent inclination or at higher ALMA resolution, another possible explanation might be angular momentum transferred from the orbit of , which could be informed by further constraining the stellar mass and properties of .
@@ Appendix A Data tables
In this section we present the following data in numerical format. TableĀ 3 lists the brightnesses, distances and periods for the Galactic supergiant stars from Fig.Ā 1. TableĀ 4 lists the phase offset values we calculated for the stars presented by Nicholls etĀ al. (2009), whereas TablesĀ 5 and 6 contains the phase offset for the subset of Gaia LSP candidates that we include in Fig.Ā 3.

### v2

@@ 4.8 Hypothesis 8: Modulation by a companion
In this section, we overview the ways in which a companion could induce an LSP (see, e.g. Wood etĀ al., 1999a, 2004b; Retter, 2005; SoszyÅski, 2007; SoszyÅski etĀ al., 2021; Percy & Shenoy, 2023, and discussions and references therein). This family of solutions is the generally favored mechanism for producing LSPs, as this class of scenarios naturally explains the long timescale and steady period as the orbital time of the companion. As discussed in SectionĀ 2, the existence of a P-L relation for LSPs is not an intrinsic problem, as orbital architecture is sensitive to the stellar mass and radius. A remaining point of tension is the relatively high fraction of long-period variables with LSPs and RV modulations compared to main-sequence stars (Nicholls etĀ al., 2009), which can be addressed by appealing to mass growth of the orbiting body (e.g. Livio & Soker 1984; see discussion in §5 of SoszyÅski etĀ al. 2021). At least regarding RSGs, the observed binary fraction is very high for massive stars (Sana etĀ al., 2012; de Mink etĀ al., 2013), and this fraction does not include undetected lower-mass companions (we return to this in SectionĀ 5.2).
@@ 4.8.3 Hypothesis 8c: Circumstellar Modulation Near Companion
| Misidentified FM | 4.1 | ā+ | X | ? | - | ? | ā+ |
| Giant convective cells | 4.2 | ā | ? | ā | - | - | X |
@@
| Non-radial pulsation | 4.6 | ā+ | ā | ? | ? | ? | X |
| Dust- | 4.7 | ā+ | ā+ | X | ā+ | ā+ | ā |
| Binarity: tidal | 4.8.1 | X | ā+ | ? | ? | ? | ā+ |
| Binarity: occultation | 4.8.2 | ā+ | ā+ | ā+ | ā+ | X | ā+ |
| Binarity: dust modulation | 4.8.3 | ā+ | ā+ | ā+ | ā+ | ā+ | ā+ |
@@
The connection between binarity, convection, pulsation, dust condensation, and mass loss has been explored by various groups in both simulations and observations (see, e.g., Höfner & Olofsson 2018 for a recent review in the context of AGB mass loss). Both 1D and 3D simulations of evolved cool giant stars exhibit dust-driven winds forming at a few times the stellar radius, seeded by (asymmetric) material which gets shocked and levitated by the pulsating stellar surface (see, e.g., Freytag & Höfner 2008; Freytag et al. 2017; Freytag & Höfner 2023; Liljegren et al. 2016, 2017, 2018; Bladh et al. 2019a, b; Siess et al. 2022; Steinwandel et al. 2022; Wiegert et al. 2024). The outflows in these works exhibit a wide range of circumstellar densities and mass-loss rates depending properties near the stellar surface, such as the pressure scale height, temperature, pulsation period, and degree of asymmetry (as well as composition; see, e.g. Karakas & Lattanzio 2014; Lattanzio & Karakas 2016).
Likewise, observations indicate a correspondence between the onset of dusty winds and different pulsation sequences (e.g. McDonald & Trabucchi, 2019; Yu etĀ al., 2021). LSPs have long been associated with dusty activity, as Wood & Nicholls (2009) found increased mid-IR emission with no difference in color for RGB stars with LSPs compared to those without, interpreted to indicate the presence of dust in a clumpy or disk-like configuration. It has also been shown in hydrodynamic simulations of AGB star winds that a companion may modify the circumstellar matter formed by dusty outflows (Maes etĀ al., 2021), with differing geometry depending on the properties of the outflow and companion.
@@
A dust-modulating companion also leaves open the possibility of secondary eclipses to be seen in the IR, which represent strong evidence for the binarity scenario in other cases (SoszyÅski etĀ al., 2021). The primary difference is that the companion would be 180 degrees out of phase with the bulk of the circumstellar dust rather than dragging the dust along as in the occultation scenario. This phase difference would potentially but not necessarily yield differences in the (highly-uncertain) expected dust geometry, depending in part on whether the dust modulation is dynamical or radiative in nature.
@@
We thus conclude that circumstellar dust modulation by a companion is the most plausible explanation for the LSP in Betelgeuse. Within this framework, we can use the observational data available to estimate the properties of a companion consistent with the RV and RV-LC phase difference. This companion is hereafter referred to as 333āBā stands for Betelbuddy. The properties (SectionĀ 5) and detectability (SectionĀ 5.2) of are discussed subsequently.
@@ 4.9 Summary of Hypotheses
TableĀ 2 provides a quick-look reference for the explanatory power and limitations of each theory discussed in this section.
@@ 5.1 Properties of Companion
where is the velocity amplitude and ād is the LSP period from the RV observations of Granzer etĀ al. (2022), and is the present-day mass estimate of from Joyce etĀ al. (2020). For simplicity, we assume a circular orbit. This assumption is further justified given the expected effects of dust reducing eccentricity, investigated primarily in the context of planets in dusty disk environments by, e.g. Li etĀ al. (2019), Coleman etĀ al. (2022). This gives us a lower mass limit of . This value is an order of magnitude larger than the mass estimates calculated for LSPs in RGBs with RV constraints, which are typically in the or less range, i.e., closer to brown dwarfs than to dwarf stars (Nicholls etĀ al., 2009). In order to cause substantial modulation over the course of an orbit, we assume that a relatively low inclination relative to our line of sight (or a high inclination relative to the plane perpendicular to our line of sight, as defined conventionally for binary stars and exoplanets) is needed for any interaction with the circumstellar dust to be observable from our vantage point. Therefore, we estimate that the mass of is unlikely to exceed , or at most with a vertically extended cloud.
@@
which recovers an orbital separation of assuming the present-day mass of Betelgeuse estimated by Joyce etĀ al. (2020). Using their stellar radius of yields an orbital distance of . This is well outside of the CO2-H2O MOLsphere described by MontargĆØs etĀ al. (2021), which extends to . It also places the L1 Lagrange point at , or . In this orbital configuration, the surface of Betelgeuse is not expected to deform significantly, as the Roche radius for this orbit is , and the equipotential corresponding to is nearly spherical. From the calculated separation, we can estimate a transit probability of , and a corresponding minimal inclination of for a grazing transit of (where an inclination of means an edge-on orbit). Therefore, even with an extended dust feature that can reach higher inclinations, our assumption of low inclination relative to the observerās line of sight is justified.
@@ 5.2 Observability of Companion
Even at its highest plausible massāsay āit would be nearly impossible to detect close to such a luminous and intrinsically variable star (the average brightness of Betelgeuse hovers near ) except in some exotic and improbable scenarios (see SectionĀ 5.2.1). Even adopting an optimistic 10 for , the luminosity contrast would still be , likely undetectable in integrated light. This challenge is exacerbated by the fact that Betelgeuseās intrinsic variability from convection and pulsations entails mag fluctuations in the V-band (see our Fig.2), with intrinsic fluctuations on the order of more than , at least 1000 times the intrinsic brightness of .
Spectral differencing at different phases would be another possible avenue for detection, but again, subtle changes in the of Betelgeuse and changes on the FM timescale would overwhelm variations in spectra unless the companion were substantially hotter than Betelgeuse. At its hottest, a 2 star might be spectral class F, corresponding to a of K. A companion of this temperature would only just barely be visible on the Wien tail in the UV (a 7000K blackbody outshines a brighter, 3600K blackbody only above Hz, or below Ć ), and realistically, this is a drastic overestimate of the temperature of .
@@ 5.2.1 Exotic Scenarios
While the companionās mass and proximity to a massive star might lead one to speculate that it may be a neutron star or other exotic compact body, as far as the authors of this study are aware, no such detections have been made (Posson-Brown etĀ al., 2006; Kashyap etĀ al., 2020). The current lack of X-ray evidence notwithstanding, there are a few reasons not to entirely ignore a neutron-star-as-companion scenario. First, it is more common than not that high-mass stars such as Betelgeuse are born as binary or even triple systems (Sana etĀ al., 2012; de Mink etĀ al., 2013; Toonen etĀ al., 2018, 2020, 2022; Offner etĀ al., 2023), many of which will produce runaway stars when the most massive star in the system explodes (e.g. Eldridge etĀ al., 2011; Zapartas etĀ al., 2021). Second, the anomalous proper motion of Betelgeuse (e.g. Harper etĀ al., 2008) still requires explanation, and a neutron star companion could provide one. In that scenario, the progenitor to a neutron star born at the same time as Betelgeuse must have been more massive than Betelgeuse in order to die earlier in a Supernova explosion. The past SN explosion of Betelgeuseās companion would have then been energetic enough to provide the necessary ākickā that sent Betelgeuse along its present observed trajectory, with the binary widening but without disrupting the system (Renzo etĀ al., 2019). Future X-ray observations scheduled at favorable epochs might rule this out conclusively. Such an experiment could also help characterize the presence of a possible corona around a non-compact companion, which might yield insight into the exact mechanism for interaction with any circumstellar dust.
@@ 5.2.2 Observability Under Normal Conditions
Another possibility is to inspect the shape of evolving asymmetries in sub-milimeter observations of the dusty region around Betelgeuse using an instrument like ALMA. There is indeed substantial asymmetry in the dust observed near the radius of the predicted orbital separation of (e.g. Richards etĀ al., 2013; OāGorman etĀ al., 2017, 2020; Kervella etĀ al., 2018b).
@@
A similar observing strategy was undertaken by OāGorman etĀ al. (2015), who took VLA measurements at 0.7, 1.3, 2.0, 3.5, and 6.1cm at multiple different phase points along the LSP over a 4-year cycle (in addition to single-epoch observations by Lim etĀ al. (1998) 4 years prior). Their Fig.Ā 1 does indicate some time-dependent modulation of an asymmetric circumstellar envelope visible in the 0.7cm interferometric data around mas444At maximum separation, will be at mas for a stellar optical photosphere diameter of 43 mas), but it is far from conclusive evidence, as an 8-year time span entails measurements during multiple different LSP cycles (orbits of ). While we do expect intrinsic diversity in dust structures on a replenishment timescale commensurate with the FM as the FM interacts with the large-scale convection, would cause a systematic modulation in the circumstellar dust (the proposed mechanism for the LSP itself).
@@ 6 Conclusions
In Section 5, we calculated physical properties of the companion and the Orionis orbital system as constrained by existing observations. A summary of these parameters is provided in TableĀ 3. We discussed why has not been detected before, commented on exotic physical scenarios that would be theoretically observable, discussed practical limitations on the companionās detectability, and described the advances in method, technique and instrumentation needed to detect it in the future.
While our work was in the refereeing process, MacLeod etĀ al. (2024) independently found astrometric evidence that a companion scenario can explain Betelgeuseās LSP, focusing their efforts on dissecting Betelgeuseās long baseline RV observations. They recovered similar orbital parameters and a slightly lower mass () due to a smaller inferred RV amplitude.
@@
The presence of dense circumstellar material is of reinvigorated interest after the recent Supernova 2023ixf, the closest core-collapse event of the decade, which showed strong signatures of early interaction in an asymmetric circumstellar medium (e.g. Hiramatsu etĀ al., 2023; Jacobson-GalĆ”n etĀ al., 2023; Hosseinzadeh etĀ al., 2023; Smith etĀ al., 2023; Teja etĀ al., 2023; Zimmerman etĀ al., 2024; Li etĀ al., 2024; Martinez etĀ al., 2024; Singh etĀ al., 2024; Moriya & Singh, 2024; Bostroem etĀ al., 2024) around a pulsating Red Supergiant progenitor (e.g. Soraisam etĀ al., 2023; Jencson etĀ al., 2023; Kilpatrick etĀ al., 2023; Xiang etĀ al., 2024). Moreover, supernova remnants are now being observed with unprecedented resolution in the era of JWST (e.g. recent work by Temim etĀ al., 2024; Milisavljevic etĀ al., 2024), and the proposed observations of Betelgeuseās surrounding environment, whether or not they support the presence of , will be valuable in constraining the expected level and geometry of circumstellar asymmetries that may imprint on the remnant morphology (Polin etĀ al., 2022; Mandal etĀ al., 2023, 2024). If the existence of is confirmed and a progenitor identified, one could also predict its own future evolution in the context of Betelgeuseās eventual explosion. Such simulations have recently been carried out in the case of Type Ia SNe with a surviving companion (Bauer etĀ al., 2019; Bhat etĀ al., 2024; Wong etĀ al., 2024).
Observations of Betelgeuse have identified photometric variations on a 30 year period consistent with a rotational km/s (Joyce etĀ al., 2020), which for an inclination angle of 20ā recovers an equatorial velocity of 15Ā km/s (see Dupree etĀ al. 1987; Gilliland & Dupree 1996; Uitenbroek etĀ al. 1998; Kervella etĀ al. 2009, 2018b, and review byWheeler & Chatzopoulos 2023). This has led to speculation that Betelgeuse may be a merger product (e.g. Wheeler etĀ al., 2017; Chatzopoulos etĀ al., 2020; Shiber etĀ al., 2024). This hypothesis was invoked to account for fast envelope rotation which would not survive the spin-down as Betelgeuse ascended up the RSG branch in a single-star evolutionary scenario. Although the large-scale convective envelope itself might be the source of this RV signal (Ma etĀ al., 2024), if the apparent RV dipole signal persists at a consistent inclination or at higher ALMA resolution, another possible explanation might be angular momentum transferred from the orbit of , which could be informed by further constraining the stellar mass and properties of .
@@ Appendix A Data tables
In this section we present the following data in numerical format. TableĀ 4 lists the brightnesses, distances and periods for the Galactic supergiant stars from Fig.Ā 1. TableĀ 5 lists the phase offset values we calculated for the stars presented by Nicholls etĀ al. (2009), whereas TablesĀ 6 and 7 contains the phase offset for the subset of Gaia LSP candidates that we include in Fig.Ā 3.